# Narrating the Video: Boosting Text-Video Retrieval via Comprehensive Utilization of Frame-Level Captions

### v1

@@ Abstract
11footnotetext: These authors contributed equally to this work.22footnotetext: Corresponding author.
@@ 1 Introduction
With the increasing popularity of various video streaming platforms, text-to-video retrieval is emerging as an important topic in both academia and industry. In order to achieve accurate retrieval results, text-video retrieval encounters two challenging points: 1) a cross-modality gap from the heterogeneous data (vision and language) and 2) time-varying information in video. To address cross-modality challenges, many studies [24, 25, 11, 9, 34, 38, 36, 22, 42] have leveraged vision-language pre-trained models such as CLIP [29] or BLIP [18] as a text and video backbone to resolve cross-modality gap between text-to-video domain. Moreover, to capture the time-varying relationships for video data, several studies [24, 22, 42, 9] have employed temporal modules, such as transformer encoder [33] or 3D convolution [32], over video frames.
@@ 2.2 Vision-Language Model with Generated Information
Figure 2: An overview of the proposed framework, NarVid. The method first generates frame-level captions (narration) for each video. Using the frame-level features of the video and narration, enhanced features are obtained through cross-modal interaction with co-attention and temporal block. (a) These enhanced features are further refined using query-aware adaptive filtering. (b) Then, the query-video and query-narration similarity matrices obtained through the multi-granularity matching are utilized for training and inference. (c) To enhance the discriminative ability of the model, we additionally use a cross-view hard negative loss during training.
@@ 3.1 Overall Process
$\boldsymbol{v^{*}}=\arg\max_{v\in V}s(\boldsymbol{q},\boldsymbol{v}).$ (1)
@@
Pipeline. The proposed framework for calculating similarity follows the process outlined in Fig. 2. A query and a video are represented as $\boldsymbol{q}=(w_{1},\dots,w_{L},w_{EOS]})$ and $\boldsymbol{v}=(f_{1},\dots,f_{K})$ using clip encoders, where $w_{L}$ is the last word of the query, $w_{EOS]}$ denotes a special token representing the query’s global meaning, and $f_{K}$ is the last frame of the video. Additionally, we generate a sequence of captions for individual frames using the narration generation module, represented as $\boldsymbol{n}=(c_{1},\dots,c_{K})$ (Sec. 3.2). The encoded features of the narration and video are enhanced through mutual co-attention and temporal transformation (Sec. 3.3). Query-aware adaptive filtering refines these features based on the query (Sec. 3.4), then query-video and query-narration matching is employed to obtain the final similarity score (Sec. 3.5). A cross-view hard negative loss is also defined by combining query-video and query-narration similarities to enhance retrieval performance (Sec. 3.6.1).
@@ 3.2 Narration Generation from Video
In this paper, we try to solve some limitations of the previous work [39] that uses a video-level caption to describe a video. Since a video-level caption is usually generated by focusing on the key scenes, it is likely to miss other narrative flows in the video clip. In addition, when a single generated caption is used to represent the entire video frames, any errors (e.g., hallucinations) in the generation model can severely affect the retrieval performance.
To address these issues, we propose the concept of narration. For each image frame $f_{k}$ in a video $\boldsymbol{v}=(f_{1},\dots,f_{K})$, we apply a large multimodal model (LMM) [21] to independently interpret it and generate a caption $c_{k}$, so that we have a sequence of captions, $\boldsymbol{n}=(c_{1},\dots,c_{K})$. Note that our framework does not focus on any specific caption generation model, the choice of captioning model and prompts can be flexible, considering various environments. The generated captions arranged in chronological order can linguistically express the semantic changes in the video over time. The generated narration is subsequently used in various ways in our proposed NarVid to improve the performance of text-to-video retrieval.
@@ 3.3 Video-Narration Cross-Modal Interaction
Each caption $c_{k}$ in the generated narration can be paired with a video frame $f_{k}$ to interact and enhance mutual information. Specifically, we use the co-attention structure [23] for interactions, a method commonly employed for cross-modal information enrichment. Fig. 2 shows the cross-modal interaction between each video $\boldsymbol{v}$ and its corresponding narration $\boldsymbol{n}$. In this context, the transformer blocks perform co-attention operations with cross-modal inputs to obtain the enhanced features $\boldsymbol{\tilde{v}}$ and $\boldsymbol{\tilde{n}}$, defined as
$(\boldsymbol{\tilde{v}},\boldsymbol{\tilde{n}})=\phi_{co-att}(\boldsymbol{v},\boldsymbol{n}).$ (2)
@@
The obtained sequential features $\boldsymbol{\tilde{v}}=(\tilde{f}_{1},\dots,\tilde{f}_{K})$ and $\boldsymbol{\tilde{n}}=(\tilde{c}_{1},\dots,\tilde{c}_{K})$ are then passed through a temporal block [24] to capture temporal relationships in the sequence of each modality, producing temporally enhanced features as follows:
$\boldsymbol{\check{v}}=\phi_{temp}(\boldsymbol{\tilde{v}}),\quad\boldsymbol{\check{n}}=\phi_{temp}(\boldsymbol{\tilde{n}}).$ (3)
@@ 3.4 Query-Aware Adaptive Filtering
To filter out query-irrelevant frames or erroneous captions in the narration, we propose a query-aware filtering module. Conventional video frame filtering methods employ a top- approach [38, 15], selecting the top- frames with the highest similarity to the query. However, using a fixed for filtering may cause a loss of important frames or the inclusion of irrelevant ones. To overcome this limitation, we adopt the nucleus sampling method [13] to implement nucleus filtering, allowing the adaptive determination of the number of keyframes.
Fig. 2 (a) illustrates our nucleus filtering process. For a given query $\boldsymbol{q}=(w_{1},\dots,w_{L},w_{EOS]})$ and the enhanced features $\boldsymbol{\check{v}}=(\check{f}_{1},\dots,\check{f}_{K})$ and $\boldsymbol{\check{n}}=(\check{c}_{1},\dots,\check{c}_{K})$, we first compute the frame-level similarities between $w_{EOS]}$ and the features $\check{f}_{k}$ and $\check{c}_{k}$ ($k=1,\dots,K$). We then apply the softmax function to normalize these similarity values to the range $[0,1]$. In the feature selection phase, we start with the feature having the highest similarity and add features one by one until the cumulative similarity exceeds a predefined threshold $p$. Through this process, we obtain a pair of final sequences of features, $\boldsymbol{z}^{v}$ for video and $\boldsymbol{z}^{n}$ for narration, defined as
@@ 3.5 Query-Video-Narration Matching
Coarse-grained matching is a common method for matching overall semantics between query and video. In addition to query-video matching, a similar query-narration matching process is conducted in NarVid. Specifically, we first apply weighted pooling to the sequence of features for the video $\boldsymbol{z}^{v}$ and narration $\boldsymbol{z}^{n}$ to obtain their coarse representations, where the weights are determined by the similarity scores from query-aware filtering module (Sec. 3.4). The $w_{EOS]}$ token is used for the coarse representation of the query $\boldsymbol{q}$. Then, the coarse-grained matching scores $s_{roase}(\boldsymbol{q},\boldsymbol{z}^{v})$ and $s_{coarse}(\boldsymbol{q},\boldsymbol{z}^{n})$ are calculated using cosine similarity between their coarse representations, effectively leveraging the filtering scores as weights in the pooling process.
@@
By averaging the coarse and fine scores, we obtain two multi-grained similarity scores, $s_{qv}$ for query-video matching and $s_{qn}$ for query-narration matching, which are used for training and inference.
@@ 3.6.1 Training Objective
Contrastive loss. For a given batch set $\boldsymbol{B}=\{(\boldsymbol{q}_{i},\boldsymbol{v}_{i},\boldsymbol{n}_{i})\}_{i=1}^{B}$, we calculate the query-video similarity matrix $\boldsymbol{S}_{qv}$ and the query-narration similarity matrix $\boldsymbol{S}_{qn}$, where each element is defined as $\boldsymbol{S}_{qv}(i,j)=s_{qv}(\boldsymbol{q}_{i},\boldsymbol{v}_{j}),\quad\boldsymbol{S}_{qn}(i,j)=s_{qn}(\boldsymbol{q}_{i},\boldsymbol{n}_{j}).$ We utilize these similarity matrices to enhance retrieval performance by considering both inter-modality (query-video) and intra-modality (query-narration) aspects. This approach modifies the InfoNCE loss [28] commonly used in text-video retrieval to magnify the discrepancy between positive and negative samples from both perspectives. The modified loss functions are defined as
@@
Cross-view hard negative loss. Hard negative samples in contrastive learning mean negative samples that closely resemble positive samples, challenging the model to distinguish between them. Although several studies [4, 30] showed that more discriminative features can be obtained if such hard negative samples are well utilized for training, they exclusively focused on either inter-modality or intra-modality when selecting hard negatives. In this paper, we propose a novel hard negative loss that leverages both different views: inter-modality and intra-modality. As shown in Fig. 2 (c) using the similarity matrices $\boldsymbol{S}_{qv}$ and $\boldsymbol{S}_{qn}$, we define two index sets of hard negatives for a given query $\boldsymbol{q}_{i}$ from the views of inter-modality and intra-modality. these sets are defined as
@@
$L_{qv}^{ard}=\frac{1}{2B}\sum_{i=1}^{B}\left(\sum_{j\in H_{i}}\Delta_{qv}(i,j)+\sum_{j\in H_{i}^{T}}\Delta_{vq}(i,j)\right),$
@@
$\Delta_{qv}(i,j)$ $=\max(0,\boldsymbol{S}_{qv}(i,i)-\boldsymbol{S}_{qv}(i,j)+\eta\lambda\sigma_{i}^{qv}),$
$\Delta_{vq}(i,j)$ $=\max(0,\boldsymbol{S}_{qv}(i,i)-\boldsymbol{S}_{qv}(j,i)+\eta\lambda\sigma_{i}^{qv}).$ (8)
@@ 4.1 Experimental Settings
Implementation details. We use the baseline settings of CLIP4Clip [24] with ViT-B/16 [7] as the visual backbone and the CLIP text encoder for query and narration encoding. For a detailed analysis of MSR-VTT, we additionally use ViT-B/32 as a visual encoder. To obtain the narration, LLaVa 1.5 7B [21] is used as the baseline generative model. We use a batch size of 64 and train for 5 epochs on three datasets, except for DiDeMo, where a batch size of 24 and 10 learning epochs are used. The threshold $p$ used in query-aware filtering is set to 0.4 for ViT-B/32 and 0.5 for ViT-B/16. All experiments are conducted on two NVIDIA RTX A6000 GPUs. Additional implementation details can be found in Sec. B.2
@@ 4.3 Ablation Study
Impact of different narration generation models. We analyze how different caption generation methods affect retrieval performance on NarVid. Specifically, as shown in Fig. 3, we compared narrations from the output of several types of captioners: a traditional image-captioning model [14], popular vision-language pre-trained models [18, 19], and powerful large multimodal models [21, 20]. The results show that the use of narration improves performance on all metrics regardless of the captioner, demonstrating that the strategy we propose works for a variety of caption generation models. Detailed experimental results about the caption generation can be found in Sec. C
@@ 5 Conclusions
  • The details of the frame-level caption generation prompts in Sec. C.
  • Additional ablation studies on the effectiveness of narration utilization and cross-view hard negative loss in Sec. D.
@@ A Multi-Granularity Matching Module
Coarse-grained matching. In Fig. 5 (a), we illustrate the coarse-grained matching process. For both video and narration, the sequence of embedding vectors for frames and captions is transformed into a single vector through weighted pooling. The cosine similarity between this pooled vector and the $w_{[EOS]}$ vector of the query is then computed to obtain the coarse matching scores $s_{coarse}(\boldsymbol{q},\boldsymbol{z}^{v})$ and $s_{coarse}(\boldsymbol{q},\boldsymbol{z}^{n})$. Here, it should be remarked that the weight values for the pooling are determined by the query-aware filtering module. In the query-aware adaptive filtering module described Sec. 3.4, we evaluate individual frames and captions based on their similarities to the given query, and obtain the final filtering scores for the selected frames $\boldsymbol{z}^{v}$ and captions $\boldsymbol{z}^{n}$ through filtering and softmax. An example of the filtering scores is shown in Fig. 2 (a) By using the filtering scores as weights for pooling, we can emphasize the most relevant frames in the query-video matching process and the most relevant captions in the query-narration matching process.
Figure 5: Process of coarse-grained and fine-grain matching. In the yellow box, we utilize the results of nucleus filtering used for weight values.
Fine-grained matching. Fig. 5 (b) shows the details of the proposed fine-granularity matching process. For query-video matching, we first compute a similarity matrix with the cosine similarity of all possible pairs of frame embeddings and word embeddings, and take frame-wise and word-wise maximum to get two vectors of maximum similarity values. The frame-wise maximum similarity vector is then used to compute the inner product with the filtering score to obtain the similarity score $s_{W2F}$. We also compute a similarity score $s_{F2W}$ using a word-wise maximum similarity vector and a weight vector from an FC layer that is trained in an end-to-end manner. Finally, the two similarity scores are summed to obtain a fine-granularity score $s_{fine}(\boldsymbol{q},\boldsymbol{z}^{v})$. For query-narration matching, the same process is performed with narration embedding vectors to obtain the score $s_{fine}(\boldsymbol{q},\boldsymbol{z}^{n})$.

### v4

@@ Abstract
11footnotetext: These authors contributed equally to this work.22footnotetext: Corresponding authors, equal leading contribution.
@@ 1 Introduction
With the increasing popularity of various video streaming platforms, text-to-video retrieval is emerging as an important topic in both academia and industry. In order to achieve accurate retrieval results, text-video retrieval encounters two challenging points: 1) a cross-modality gap from the heterogeneous data (vision and language) and 2) time-varying information in video. To address cross-modality challenges, many studies [25, 26, 11, 9, 35, 39, 37, 23, 43, 45, 20] have leveraged vision-language pre-trained models such as CLIP [30] or BLIP [18] as a text and video backbone to resolve cross-modality gap between text-to-video domain. Moreover, to capture the time-varying relationships for video data, several studies [25, 23, 43, 9] have employed temporal modules, such as transformer encoder [34] or 3D convolution [33], over video frames.
@@ 2.2 Vision-Language Model with Generated Information
Figure 2: An overview of the proposed framework, NarVid. The method first generates frame-level captions (narration) for each video. (a) Using the frame-level features of the video and narration, enhanced features are obtained through cross-modal interaction with co-attention and temporal block. (b) These enhanced features are further refined using query-aware adaptive filtering. (c) Then, the query-video and query-narration similarity matrices obtained through the multi-granularity matching are utilized for training and inference. (d) To enhance the discriminative ability of the model, we additionally use a cross-view hard negative loss during training.
@@ 3.1 Overall Process
$\boldsymbol{v^{*}}=\arg\max_{\boldsymbol{v}\in V}s(\boldsymbol{q},\boldsymbol{v}).$ (1)
@@
Pipeline. The proposed framework for calculating similarity follows the process outlined in Fig. 2. A query and a video are represented as $\boldsymbol{q}=(w_{1},\dots,w_{L},w_{[EOS]})$ and $\boldsymbol{v}=(f_{1},\dots,f_{K})$ using clip encoders, where $w_{L}$ is the last word of the query, $w_{[EOS]}$ denotes a special token representing the query’s global meaning, and $f_{K}$ is the last frame of the video. Additionally, we generate a sequence of captions for individual frames using the narration generation module, represented as $\boldsymbol{n}=(c_{1},\dots,c_{K})$ (Sec. 3.2). The encoded features of the narration and video are enhanced through mutual co-attention and temporal transformation (Sec. 3.3). Query-aware adaptive filtering refines these features based on the query (Sec. 3.4), then query-video and query-narration matching is employed to obtain the final similarity score (Sec. 3.5). A cross-view hard negative loss is also defined by combining query-video and query-narration similarities to enhance retrieval performance (Sec. 3.6.1).
@@ 3.2 Narration Generation from Video
In this paper, we try to solve some limitations of the previous work [40] that uses a video-level caption to describe a video. Since a video-level caption is usually generated by focusing on the key scenes, it is likely to miss other narrative flows in the video clip. In addition, when a single generated caption is used to represent the entire video frames, any errors (e.g., hallucinations) in the generation model can severely affect the retrieval performance. To address these issues, we propose the concept of narration. For each image frame $f_{k}$ in a video $\boldsymbol{v}=(f_{1},\dots,f_{K})$, we apply a large multimodal model (LMM) [22] to independently interpret it and generate a caption $c_{k}$, so that we have a sequence of captions, $\boldsymbol{n}=(c_{1},\dots,c_{K})$. Note that our framework does not focus on any specific caption generation model, the choice of captioning model and prompts can be flexible, considering various environments. The generated captions arranged in chronological order can linguistically express the semantic changes in the video over time. The generated narration is subsequently used in various ways in our proposed NarVid to improve the performance of text-to-video retrieval.
@@ 3.3 Video-Narration Cross-Modal Interaction
Each caption $c_{k}$ in the generated narration can be paired with a video frame $f_{k}$ to interact and enhance mutual information. Specifically, we use the co-attention structure [24] for interactions, a method commonly employed for cross-modal information enrichment. Fig. 2 (a) shows the cross-modal interaction between each video $\boldsymbol{v}$ and its corresponding narration $\boldsymbol{n}$. In this context, the transformer blocks perform co-attention operations with cross-modal inputs to obtain the enhanced features $\boldsymbol{\hat{v}}$ and $\boldsymbol{\hat{n}}$, defined as
$(\boldsymbol{\hat{v}},\boldsymbol{\hat{n}})=\phi_{co-att}(\boldsymbol{v},\boldsymbol{n}).$ (2)
@@
The obtained sequential features $\boldsymbol{\hat{v}}=(\hat{f}_{1},\dots,\hat{f}_{K})$ and $\boldsymbol{\hat{n}}=(\hat{c}_{1},\dots,\hat{c}_{K})$ are then passed through a temporal block [25] to capture temporal relationships in the sequence of each modality, producing temporally enhanced features as follows:
$\boldsymbol{\check{v}}=\phi_{temp}(\boldsymbol{\hat{v}}),\quad\boldsymbol{\check{n}}=\phi_{temp}(\boldsymbol{\hat{n}}).$ (3)
@@ 3.4 Query-Aware Adaptive Filtering
To filter out query-irrelevant frames or erroneous captions in the narration, we propose a query-aware filtering module. Conventional video frame filtering methods employ a top-k approach [39, 15], selecting the top-k frames with the highest similarity to the query. However, using a fixed for filtering may cause a loss of important frames or the inclusion of irrelevant ones. To overcome this limitation, we adopt the nucleus sampling method [13] to implement nucleus filtering, allowing the adaptive determination of the number of keyframes.
Fig. 2 (b) illustrates our nucleus filtering process. For a given query $\boldsymbol{q}=(w_{1},\dots,w_{L},w_{[EOS]})$ and the enhanced features $\boldsymbol{\check{v}}=(\check{f}_{1},\dots,\check{f}_{K})$ and $\boldsymbol{\check{n}}=(\check{c}_{1},\dots,\check{c}_{K})$, we first compute the frame-level similarities between $w_{[EOS]}$ and the features $\check{f}_{k}$ and $\check{c}_{k}$ ($k=1,\dots,K$). We then apply the softmax function to normalize these similarity values to the range $[0,1]$. In the feature selection phase, we start with the feature having the highest similarity and add features one by one until the cumulative similarity exceeds a predefined threshold $p$. Through this process, we obtain a pair of final sequences of features, $\boldsymbol{z}^{v}$ for video and $\boldsymbol{z}^{n}$ for narration, defined as
@@ 3.5 Query-Video-Narration Matching
Coarse-grained matching is a common method for matching overall semantics between query and video. In addition to query-video matching, a similar query-narration matching process is conducted in NarVid. Specifically, we first apply weighted pooling to the sequence of features for the video $\boldsymbol{z}^{v}$ and narration $\boldsymbol{z}^{n}$ to obtain their coarse representations, where the weights are determined by the similarity scores from query-aware filtering module (Sec. 3.4). The $w_{[EOS]}$ token is used for the coarse representation of the query $\boldsymbol{q}$. Then, the coarse-grained matching scores $s_{coarse}(\boldsymbol{q},\boldsymbol{z}^{v})$ and $s_{coarse}(\boldsymbol{q},\boldsymbol{z}^{n})$ are calculated using cosine similarity between their coarse representations, effectively leveraging the filtering scores as weights in the pooling process.
@@
By averaging the coarse and fine scores, we obtain two multi-grained similarity scores, $s_{qv}$ for query-video matching and $s_{qn}$ for query-narration matching, which are used for training and inference. For more detailed information is qiven in Supplementary Sec. A
@@ 3.6.1 Training Objective
Contrastive loss. For a given batch set $\{(\boldsymbol{q}_{i},\boldsymbol{v}_{i},\boldsymbol{n}_{i})\}_{i=1}^{B}$, we calculate the query-video similarity matrix $\boldsymbol{S}_{qv}$ and the query-narration similarity matrix $\boldsymbol{S}_{qn}$, where each element is defined as $\boldsymbol{S}_{qv}(i,j)=s_{qv}(\boldsymbol{q}_{i},\boldsymbol{v}_{j}),\quad\boldsymbol{S}_{qn}(i,j)=s_{qn}(\boldsymbol{q}_{i},\boldsymbol{n}_{j}).$ We utilize these similarity matrices to enhance retrieval performance by considering both inter-modality (query-video) and intra-modality (query-narration) aspects. This approach modifies the InfoNCE loss [29] commonly used in text-video retrieval to magnify the discrepancy between positive and negative samples from both perspectives. The modified loss functions are defined as
@@
Cross-view hard negative loss. Hard negative samples in contrastive learning mean negative samples that closely resemble positive samples, challenging the model to distinguish between them. Although several studies [4, 31] showed that more discriminative features can be obtained if such hard negative samples are well utilized for training, they exclusively focused on either inter-modality or intra-modality when selecting hard negatives. In this paper, we propose a novel hard negative loss that leverages both different views: inter-modality and intra-modality. As shown in Fig. 2 (d) using the similarity matrices $\boldsymbol{S}_{qv}$ and $\boldsymbol{S}_{qn}$, we define two index sets of hard negatives for a given query $\boldsymbol{q}_{i}$ from the views of inter-modality and intra-modality. these sets are defined as
@@
$L_{qv}^{hard}=\frac{1}{2B}\sum_{i=1}^{B}\left(\sum_{j\in H_{i}}\Delta_{qv}(i,j)+\sum_{j\in H_{i}^{T}}\Delta_{vq}(i,j)\right),$
@@
$\Delta_{qv}(i,j)$ $=\max(0,\boldsymbol{S}_{qv}(i,j)-\boldsymbol{S}_{qv}(i,i)+\eta\lambda\sigma_{i}^{qv}),$
$\Delta_{vq}(i,j)$ $=\max(0,\boldsymbol{S}_{qv}(j,i)-\boldsymbol{S}_{qv}(i,i)+\eta\lambda\sigma_{i}^{vq}).$ (8)
@@ 4.1 Experimental Settings
Implementation details. We use the baseline settings of CLIP4Clip [25] with ViT-B/16 [7] as the visual backbone and the CLIP text encoder for query and narration encoding. For a detailed analysis of MSR-VTT, we additionally use ViT-B/32 as a visual encoder. To obtain the narration, LLaVa 1.5 7B [22] is used as the baseline generative model. We use a batch size of 64 and train for 5 epochs on three datasets, except for DiDeMo, where a batch size of 24 and 10 learning epochs are used. The threshold $p$ used in query-aware filtering is set to 0.4 for ViT-B/32 and 0.5 for ViT-B/16. All experiments are conducted on two NVIDIA RTX A6000 GPUs. Additional implementation details can be found in Supplementary Sec. B.2
@@ 4.3 Ablation Study
Impact of different narration generation models. We analyze how different caption generation methods affect retrieval performance on NarVid. Specifically, as shown in Fig. 3, we compared narrations from the output of several types of captioners: a traditional image-captioning model [14], popular vision-language pre-trained models [18, 19], and powerful large multimodal models [22, 21]. The results show that the use of narration improves performance on all metrics regardless of the captioner, demonstrating that the strategy we propose works for a variety of caption generation models. Interestingly, we observed that more powerful VLMs generate detailed captions containing many query-irrelevant words. These well-generated but query-misaligned captions could negatively impact retrieval performance. Detailed experimental results about the caption generation can be found in Supplementary Sec. C
@@ 5 Conclusions
Acknowledgments
This work was supported in part by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (No. RS-2021-II212068, Artificial Intelligence Innovation Hub) in part by the MSIT(Ministry of Science and ICT), Korea, under the ITRC(Information Technology Research Center) support program(IITP-2024-RS-2024-00437718) supervised by the IITP(Institute for Information & Communications Technology Planning & Evaluation), and in part by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (No.RS-2023-00227431, Development of 3D space digital media standard technology).
@@ Acknowledgments
  • The details of the frame-level caption generation process in Sec. C.
  • Additional ablation studies in Sec. D.
@@ A Multi-Granularity Matching Module
Coarse-grained matching. In Fig. 5 (a), we illustrate the coarse-grained matching process. For both video and narration, the sequence of embedding vectors for frames and captions is transformed into a single vector through weighted pooling. The cosine similarity between this pooled vector and the $w_{[EOS]}$ vector of the query is then computed to obtain the coarse matching scores $s_{coarse}(\boldsymbol{q},\boldsymbol{z}^{v})$ and $s_{coarse}(\boldsymbol{q},\boldsymbol{z}^{n})$. Here, it should be remarked that the weight values for the pooling are determined by the query-aware filtering module. In the query-aware adaptive filtering module described in Sec. 3.4, we evaluate individual frames and captions based on their similarities to the given query, and obtain the final filtering scores for the selected frames $\boldsymbol{z}^{v}$ and captions $\boldsymbol{z}^{n}$ through filtering and softmax. An example of the filtering scores is shown in Fig. 2 (b)
By using the filtering scores as weights for pooling, we can emphasize the most relevant frames in the query-video matching process and the most relevant captions in the query-narration matching process.
Figure 5: Process of coarse-grained and fine-grain matching. In the yellow box, we utilize the results of nucleus filtering for weight values.
Fine-grained matching. Fig. 5 (b) shows the details of the proposed fine-granularity matching process. For query-video matching, we first compute a similarity matrix with the cosine similarity of all possible pairs of frame embeddings and word embeddings and take frame-wise and word-wise maximums to get two vectors of maximum similarity values. The frame-wise maximum similarity vector is then used to compute the inner product with the filtering score to obtain the similarity score $s_{W2F}$. We also compute a similarity score $s_{F2W}$ using a word-wise maximum similarity vector and a weight vector from an FC layer that is trained in an end-to-end manner. Finally, the two similarity scores are summed to obtain a fine-granularity score $s_{fine}(\boldsymbol{q},\boldsymbol{z}^{v})$. For query-narration matching, the same process is performed with narration embedding vectors to obtain the score $s_{fine}(\boldsymbol{q},\boldsymbol{z}^{n})$.